OpenStudy (anonymous):

In the equation x/4 + x = 6, x = 1.2 true or false

OpenStudy (anonymous):

x/4 + x = 6 x/4 + x(4/4) = 6 x/4 + 4x/4 = 6 5x/4 = 6 x(5/4) = 6 x(1.2) = 6 x = 6/(1.2) x = 24/5 False

OpenStudy (anonymous):

if u put x=1.2=(6/5) into (x/4)+x the asnwer is 1.5 \[1.5 \neq 6\] so; the answer is FALSE